Chickentender37:

Last one :o A passenger plane left New York four hours before a cargo plane. The planes flew in opposite directions. The cargo plane flew at 400 mph for seven hours. After this time the planes were 5220 mi. apart. What was the passenger plane's speed?

Vocaloid:

 After the cargo plane has been travelling for 7 hours, they are 5220 miles apart. At that instant, the cargo plane has travelled 7*400 = 2800 miles Meaning the passenger plane has travelled 5220 - 2800 = 2420 miles at that point Now, remember, the passenger plane got a 4 hour head start, so the passenger plane has been travelling 7 + 4 = 11 hours at this point. passenger plane speed = distance/time = 2420 miles/11 hours